OpenStudy (anonymous):

Line AB has the equation 3x - 4y + 5 = 0. Point (p, p+2) lies on line AB. What is the value of p?

OpenStudy (anonymous):

@satellite73

OpenStudy (mayankdevnani):

just plug ( p ,p+2) in the equation of line and solve for p !

OpenStudy (anonymous):

What would be your answer?

OpenStudy (mayankdevnani):

first solve it your own and then show your work and i'll tell you are you correct or wrong ?

OpenStudy (anonymous):

-3 was my answer :)

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

yup ! you are correct ! good job :)

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

check my work

OpenStudy (mayankdevnani):

jost post it !

OpenStudy (anonymous):

What is the gradient of line AB from the previous question?

OpenStudy (anonymous):

I said 3/4

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

you are correct !

OpenStudy (anonymous):

I am??

OpenStudy (mayankdevnani):

yeah

OpenStudy (anonymous):

explain how u did it

OpenStudy (anonymous):

I said : You need to rearrange in to the form y=mx+c, where m is the gradient and c is the y intercept. In this case the rearranges equation is y=(3/4)x+(5/4)

OpenStudy (mayankdevnani):

\[\large \bf ax+by+c=0\] let this be a equation of line. formula for finding slope of this line :- \[\large \bf slope=\frac{-a}{b}\]

OpenStudy (mayankdevnani):

just plug the values in it and solve !

OpenStudy (anonymous):

A geometric series has first term 80 and common ratio 1/2. What are the next two terms?

OpenStudy (anonymous):

I said 40 and 20

OpenStudy (mayankdevnani):

show your work !

OpenStudy (anonymous):

Multiply the first term by 1/2 to get the second term (40) and do it again to get the third (20).

OpenStudy (mayankdevnani):

correct !

OpenStudy (anonymous):

What would you use to find the approximate area of an integral between certain limits?

OpenStudy (anonymous):

@mayankdevnani ?

OpenStudy (mayankdevnani):

i don't know much about this !

OpenStudy (anonymous):

I said the Trapezium rule

OpenStudy (mayankdevnani):

sorry !

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

What grade u in?

OpenStudy (mayankdevnani):

11 grade !

OpenStudy (mayankdevnani):

and wbu ?

OpenStudy (anonymous):

11

OpenStudy (mayankdevnani):

great !